OpenStudy (anonymous):

Simplify open parentheses x to the 4 fifths power close parentheses to the 3 sevenths power

OpenStudy (anonymous):

\[\huge (x^5)^{\frac{3}{7}}\]?

OpenStudy (anonymous):

\[\huge (x^\frac{4}{5})^{\frac{3}{7}}\]

OpenStudy (anonymous):

yes it looks like that

OpenStudy (anonymous):

multiply the exponents

OpenStudy (anonymous):

i.e. compute \[\frac{4}{5}\times \frac{3}{7}\]

OpenStudy (anonymous):

so then you would get \[x \frac{ 12 }{ 35 }\]

OpenStudy (anonymous):

yes, with that as the exponent, right

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw